the length of a rectangle is four more than five it's width. Its perimeter is 56 inches. Find its dimensions (length and width).

Answer by ewatrrr(24785)

 
Hi
.
l = 5w + 4
(5w+4) + w = 56/2
6w = 28-4
6w = 24
w = 4 and l = 24